OpenStudy (anonymous):

A gardener weeds 2/3 of a flower bed in 6 minutes. How much of the flower bed did he weed in 1 minute? ?/?

OpenStudy (anonymous):

6 / 1 = ?

OpenStudy (anonymous):

@iGreen. @S.Marie23 @sleepyjess @Rainbow_dashie

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No, check again.

OpenStudy (littlebird):

There are multiple ways you can solve this: (2/3)/6=2/18=1/9 or 2/3 in 6 minutes I find how long it takes to finish a third which is 6/2=3 minutes I multiply 3 min by 3 to find how long it takes to finish 3/3 of the garden This tells me that 1 garden=9 minutes How much garden is done in 1 minute is equal to 1/9